OpenStudy (anonymous):

solve the system by the substitution method 18x-4y=4

OpenStudy (anonymous):

18x-4y=4 pretend that the y value is not there 18x=4 x=4.5 18(4.5)-4y=4 solve for y 81-4y=4 -4y=-77 y=19.25 Don't know if this is correct....